OpenStudy (anonymous):

Find the distance between the points (-2, 4) and (4, -6).

OpenStudy (solomonzelman):

Let me give you the distance formula. \(\Large\color{black}{ \displaystyle {\rm D} =\sqrt{ (\color{blue}{{\rm y}_1}-\color{red}{{\rm y}_2})^2+(\color{green}{{\rm x}_1}-\color{darkgoldenrod}{{\rm x}_2})^2 }}\) where \(\Large\color{black}{ \displaystyle {\rm D} }\) is the distance. \(\Large\color{black}{ \displaystyle (\color{green}{{\rm x}_1}~,~~\color{blue}{{\rm y}_1}) }\) and \(\Large\color{black}{ \displaystyle (\color{darkgoldenrod}{{\rm x}_2}~,~~\color{red}{{\rm y}_2}) }\) are your two points. ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~ IN YOUR CASE, the points are \(\Large\color{black}{ \displaystyle (\color{green}{-2}~,~~\color{blue}{4}) }\) and \(\Large\color{black}{ \displaystyle (\color{darkgoldenrod}{4}~,~~\color{red}{-6}) }\).

OpenStudy (anonymous):

Would the answer be D=136